# The Complexity of Minimum-Envy House Allocation Over Graphs

### v2

@@ 4.2 Balanced Separator Based Algorithm
For example, the class of planar graphs are $\mathcal{O}(\sqrt{n})$-balanced-separable (due to the planar separator theorem [14]), graphs of treewidth tw are $(\text{tw}+1)$-balanced-separable.
@@
Annnotated Optimal House Allocation of Agent Network Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$. Question: Compute an allocation $\phi$ satisfying $\phi(a)\in\mathcal{F}_{a}$ which minimizes |$|\{a\in\mathcal{A}\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in\mathcal{A},\phi(a^{\prime})\in\mathcal{P}_{a}\}|+|\{a\in B\mid\phi(a)\notin\mathcal{P}_{a}\}|$
@@
1:procedure OHAANR($\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$)
@@
3:         return 00 $\triangleright$ No agent is envious
@@
5:     $\xi\leftarrow\infty$triangleright In the worst case, no assignment exists respecting $(\mathcal{F}_{a})_{a\in\mathcal{A}}$
@@
7:     $\mathcal{A}_{1},\mathcal{A}_{2}\leftarrow$ the separated parts of $\mathcal{A}\setminus S$triangleright $|\mathcal{A}_{1}|,|\mathcal{A}_{2}|\leqslant 2n/3$
@@
9:         $C\leftarrow\{a\in S\mid\phi(a)\in\mathcal{P}_{a}\}$triangleright Agents with preferred house
10:         $D\leftarrow\{a\in S\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$triangleright Envious, non-angry
11:         $Q\leftarrow\{a\in S\cap B\mid a\notin\mathcal{P}_{a}\}$triangleright Envious, angry agents
12:         $R\leftarrow S\setminus(C\cup D\cup Q)$triangleright Other agents
13:         $B^{\prime}\leftarrow B\cup\{a^{\prime}\in\mathcal{A}\setminus S\mid\exists a\in S,\{a,a^{\prime}\}\in E_{\mathcal{A}},\phi(a)\in\mathcal{P}_{a^{\prime}}\}$triangleright Updated angry agents
@@
[…]
@@
$=2^{\mathcal{O}\left(\log n\cdot\left(\frac{n}{\log n}+\frac{2n/3}{(3/4)\log(n)}+\frac{4n/9}{(9/16)\log(n)}+\cdots\right)+n\right)}$ as ($\text{as }(3/4)^{r}\log n\leqslant\log((2/3)^{r}n)$
@@
In particular, these include planar graphs, bounded genus graphs [8], and any proper minor closed graph classes [13].
@@
[…]
@@ 5 Complexities for various Vertex Cover sizes of the Agent Graph
    For some guess of $C$, the total number of envious agents in $S$ is at most $(k-|C|)$, while the total number of envious agents in $\mathcal{A}\setminus S$ is $\zeta$. Since we are iterating over all guesses of $C$, for any minimum-envy allocation, there would exist a guess $C$ such that $(k-|C|)$ is exactly the number of envious agents. Hence, to solve Optimal House Allocation of Agent Network, it suffices to output the minimum of $k-|C|+\zeta$ over all guesses.
@@
1:$\xi\leftarrow n$triangleright In the worst case, all agents are envious
@@
3:     $J\leftarrow\{a\in S\mid\phi(a)\notin\mathcal{P}_{a}\text{ and }\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$triangleright already envious
@@
This motivates us to explore the exact boundary of tractable and intractable cases for various sizes of the minimum vertex cover. A natural question would be to ask whether vertex covers of size at most $n^{\varepsilon}$ for any constant $\varepsilon\in(0,1)$ would also allow efficient algorithms. We show in the next couple of theorems, that this is indeed not the case; Optimal House Allocation of Agent Network in $\mathsf{NP}$-hard even for bipartite graphs and split graphs with a vertex cover size of at most $n^{\varepsilon}$ for every constant $\varepsilon\in(0,1)$. Inspired from the reductions due to Madathil et. al. [15], the reductions we provide work even when each house is preferred by constantly many agents.
@@
    Let =$\mathcal{H}=\{h_{e}\mid e\in E\}\cup\{h^{*}_{j}\mid j\in[N+2M-\binom{k}{2}]\}$. Therefore, $|\mathcal{H}|=M+N+2M-\binom{k}{2}=|\mathcal{A}|+(M-\binom{k}{2})$.
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $N+2M-(N+2M-\binom{k}{2})=\binom{k}{2}$ many such agents, each of the form $a^{1}_{e}$ for some $e\in E$. Let $\gamma=\binom{k}{2}$. Let $a^{1}_{e_{1}},a^{1}_{e_{2}},\ldots,a^{1}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. For any $v\in V$, $v$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. Further, the number of envious agents, i.e. the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore indeed, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.
@@
    Let =$\mathcal{H}=\{h^{j}_{e}\mid e\in E,j\in[t]\}\cup\{h^{*}_{j}\mid j\in[(M-\binom{k}{2})t+N]\}$. Therefore, $|\mathcal{H}|=M\cdot t+N+(M-\binom{k}{2})t=|\mathcal{A}|+(M-\binom{k}{2})t$.
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $Mt-(M-\binom{k}{2})t=\binom{k}{2}t$ many such agents. Let $\gamma=\binom{k}{2}t$. Let $a^{j_{1}}_{e_{1}},a^{j_{2}}_{e_{2}},\ldots,a^{j_{\gamma}}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. Therefore, the number of distinct edges in $e_{1},e_{2},\ldots,e_{\gamma}$ is at least $\gamma/t=\binom{k}{2}$. Moreover, for any $v\in V$, $v$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. However, the number of envious agents, i.e., the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.

### v3

@@ 4.2 Balanced Separator Based Algorithm
For example, the class of planar graphs are $\mathcal{O}(\sqrt{n})$-balanced-separable (due to the planar separator theorem [lipton1979separator]), graphs of treewidth tw are $(\text{tw}+1)$-balanced-separable.
@@
Annnotated Optimal House Allocation of Agent Network Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$. Question: Compute an allocation $\phi$ satisfying $\phi(a)\in\mathcal{F}_{a}$ which minimizes $|\{a\in\mathcal{A}\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in\mathcal{A},\phi(a^{\prime})\in\mathcal{P}_{a}\}|+|\{a\in B\mid\phi(a)\notin\mathcal{P}_{a}\}|$
@@
[…]
@@
3:   return 0 $\triangleright$ No agent is envious
@@
5:  $\xi\leftarrow\infty$ $\triangleright$ In the worst case, no assignment exists respecting $(\mathcal{F}_{a})_{a\in\mathcal{A}}$
@@
7:  $\mathcal{A}_{1},\mathcal{A}_{2}\leftarrow$ the separated parts of $\mathcal{A}\setminus S$ $\triangleright$ $|\mathcal{A}_{1}|,|\mathcal{A}_{2}|\leqslant 2n/3$
@@
9:   $C\leftarrow\{a\in S\mid\phi(a)\in\mathcal{P}_{a}\}$ $\triangleright$ Agents with preferred house
10:   $D\leftarrow\{a\in S\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$ $\triangleright$ Envious, non-angry
11:   $Q\leftarrow\{a\in S\cap B\mid a\notin\mathcal{P}_{a}\}$ $\triangleright$ Envious, angry agents
12:   $R\leftarrow S\setminus(C\cup D\cup Q)$ $\triangleright$ Other agents
13:   $B^{\prime}\leftarrow B\cup\{a^{\prime}\in\mathcal{A}\setminus S\mid\exists a\in S,\{a,a^{\prime}\}\in E_{\mathcal{A}},\phi(a)\in\mathcal{P}_{a^{\prime}}\}$ $\triangleright$ Updated angry agents
@@
Now, consider inputs with the guarantee $m=\mathcal{O}(n)$. For treewidth $\leqslant\text{tw}$ graphs, this algorithms has worst case running time of $2^{\mathcal{O}(\text{tw}\log^{2}n+n)}$. For $f(n)=\mathcal{O}\left(\frac{n}{\log n}\right)$, the algorithm has the following worst case runtime:
@@
$=2^{\mathcal{O}\left(\log n\cdot\left(\frac{n}{\log n}+\frac{2n/3}{(3/4)\log(n)}+\frac{4n/9}{(9/16)\log(n)}+\cdots\right)+n\right)}$ $\text{as }(3/4)^{r}\log n\leqslant\log((2/3)^{r}n)$
@@
In particular, these include planar graphs, bounded genus graphs [gilbert1984separator], and any proper minor closed graph classes [kawarabayashi2010separator].
@@
    $2^{\mathcal{O}(\text{tw}\log^{2}n+n)}$ for treewidth $\leqslant\text{tw}$ graphs.
@@ 5 Complexities for various Vertex Cover sizes of the Agent Graph
    For some guess of $C$, the total number of envious agents in $S$ is at most (k|C|)(k-|C|), while the total number of envious agents in $\mathcal{A}\setminus S$ is $\zeta$. Since we are iterating over all guesses of $C$, for any minimum-envy allocation, there would exist a guess $C$ such that (k|C|)(k-|C|) is exactly the number of envious agents. Hence, to solve Optimal House Allocation of Agent Network, it suffices to output the minimum of $k-|C|+\zeta$ over all guesses.
@@
[…]
@@
3:  $J\leftarrow\{a\in S\mid\phi(a)\notin\mathcal{P}_{a}\text{ and }\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$ $\triangleright$ already envious
@@
This motivates us to explore the exact boundary of tractable and intractable cases for various sizes of the minimum vertex cover. A natural question would be to ask whether vertex covers of size at most $n^{\varepsilon}$ for any constant $\varepsilon\in(0,1)$ would also allow efficient algorithms. We show in the next couple of theorems, that this is indeed not the case; Optimal House Allocation of Agent Network in $\mathsf{NP}$-hard even for bipartite graphs and split graphs with a vertex cover size of at most $n^{\varepsilon}$ for every constant $\varepsilon\in(0,1)$. Inspired from the reductions due to Madathil et. al. [madathil2024cost], the reductions we provide work even when each house is preferred by constantly many agents.
@@
    Let $\mathcal{H}=\{h_{e}\mid e\in E\}\cup\{h^{*}_{j}\mid j\in[N+2M-\binom{k}{2}$]\}. Therefore, $|\mathcal{H}|=M+N+2M-\binom{k}{2}=|\mathcal{A}|+(M-\binom{k}{2})$.
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $N+2M-(N+2M-\binom{k}{2})=\binom{k}{2}$ many such agents, each of the form $a^{1}_{e}$ for some $e\in E$. Let $\gamma=\binom{k}{2}$. Let $a^{1}_{e_{1}},a^{1}_{e_{2}},\ldots,a^{1}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. For any $\in V, vv$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. Further, the number of envious agents, i.e. the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore indeed, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.
@@
    Let $\mathcal{H}=\{h^{j}_{e}\mid e\in E,j\in[t]\}\cup\{h^{*}_{j}\mid j\in[(M-\binom{k}{2})t+N]\}$. Therefore, $|\mathcal{H}|=M\cdot t+N+(M-\binom{k}{2})t=|\mathcal{A}|+(M-\binom{k}{2})t$.
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $Mt-(M-\binom{k}{2})t=\binom{k}{2}t$ many such agents. Let $\gamma=\binom{k}{2}t$. Let $a^{j_{1}}_{e_{1}},a^{j_{2}}_{e_{2}},\ldots,a^{j_{\gamma}}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. Therefore, the number of distinct edges in $e_{1},e_{2},\ldots,e_{\gamma}$ is at least $\gamma/t=\binom{k}{2}$. Moreover, for any $\in V, vv$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. However, the number of envious agents, i.e., the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.